Two cars are driving at the same constant speed on a straight road, with car 1 in front of car 2. Car 1 suddenly starts to brake with constant acceleration and stops in 10 m. At the instant car 1 comes to a stop, car 2 begins to brake with the same acceleration. It comes to a halt just as it reaches the back of car 1. What was the separation between the cars before they starting braking?
Answer #1

s peal of bothors are some as (Let) Accelerabon of botu or are alto sawe , tet as a final velrerty =V-0 have llas) distance=d- V+at at tho time of bra of as 1 2 So, seterafmofons atthe tine、 ofbaking